the length of a rectangle is three inches more then its width. the perimeter of the rectangle is 30 inches.
A. define a varibale for the width.
B.write an expression for the length in terms of the width.
C. write an equation to find the width of the rectangle. solve your equation.
D. what is the lenght of the rectangle.
please help

Answer by ewatrrr(24785)


Hi,
the length of a rectangle is three inches more then its width. the perimeter of the rectangle is 30 inches.
A.define a variable for the width. let w represent the width
B.write an expression for the length in terms of the width. the length would be (w+3)
C.write an equation to find the width of the rectangle. 2w +2(w+3) = 30
solve your equation.
2w +2(w+3) = 30
4w + 6 = 30
4w = 24
w = 6
D. what is the length of the rectangle? 6in + 3in = 9in
check your answer
2*6in + 2*9in = 30in